OpenStudy (anonymous):

5t-2r=25 solve for t HURRY

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@brucebaner @ganeshie8

OpenStudy (anonymous):

help?

ganeshie8 (ganeshie8):

\[\large 5\color{red}{t}-2r = 25\]

OpenStudy (anonymous):

@Tuta

OpenStudy (anonymous):

r= (25-5t)/2

OpenStudy (anonymous):

Can you show me how to solve for t?

ganeshie8 (ganeshie8):

\[\large 5\color{red}{t}-2r = 25\] to solve \(\large \color{red}{t}\), you need to isolate it

ganeshie8 (ganeshie8):

start by adding \(\large 2r\) to both sides what do you get ?

ganeshie8 (ganeshie8):

\[\large 5\color{red}{t} = 25+2r\]

ganeshie8 (ganeshie8):

next divide 5 both sides

ganeshie8 (ganeshie8):

\[\large \color{red}{t} = \dfrac{25+2r}{5}\]

ganeshie8 (ganeshie8):

we are done since \(\large \color{red}{t}\) is all by itself now

OpenStudy (anonymous):

Ty